Marigold Corp. manufactures widgets. Bowden Company has approached Marigold with a proposal to sell the company widgets at a price of $62320 for 100000 units. Marigold is currently making these components in its own factory. The following costs are associated with this part of the process when 100000 units are produced:

Direct material $ 23560
Direct labor 22040
Manufacturing overhead 30400
Total

$76000


The manufacturing overhead consists of $12160 of costs that will be eliminated if the components are no longer produced by Marigold. From Marigold’s point of view, how much is the incremental cost or savings if the widgets are bought instead of made?

$4560 incremental cost

$1520 incremental savings

$13680 incremental cost

$13680 incremental savings

Please try and explain the calculations!

thankssss
Answer #1

Answer: $ 4,560 Incremental cost

Cost to Make the 100,000 Units = 76,000

Cost incurred to Buying the Units = Price + Unavoidable MOH = 62320+(30400-12,160) = 80,560

Loss on Bought the units = 80,560- 76,000 = 4,560 Incremental Cost